# Numerical simulation of type II primordial black hole formation

### v5

@@ Abstract
This study investigates the formation of primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe. We find that, for a sufficiently large initial amplitude, the configuration of trapping horizons shows characteristic structure due to the existence of bifurcating trapping horizons. We call this structure of the trapping horizons “Type II PBH”, while the structure without a bifurcating trapping horizon “Type I PBH”, which is typically generated from a relatively small amplitude of the initial fluctuation. In Ref. [1], in the dust-dominated universe, the Type II PBH can be realized by the Type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case with a monotonic areal radius (Type I fluctuation). Our research reveals that a type II fluctuation does not necessarily result in a type II PBH in the case of the radiation fluid. We also find that for the initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude depending on its specific profile rather than its fluctuation type.
@@ 1 Introduction
Primordial black holes (PBHs) play a crucial role in exploring the early universe, particularly in investigating inhomogeneities whose scales are much smaller than those of the fluctuations probed by the cosmological microwave background. Since PBHs are remnants of non-linear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity. They are considered as one of the candidates for dark matter and can be detected as black hole (BH) binaries by gravitational wave interferometers. According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [2, 3, 4, 5, 6].
@@
The mechanism of PBH formation can be briefly understood by considering a model of the gravitational collapse in the flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe (see, e.g., Refs. [7, 8]). A criterion for the collapse has been firstly proposed by Carr [9] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively. A more accurate threshold value has been explored by numerically analyzing general relativistic non-linear dynamics [10, 11, 12, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$ which represents the growing adiabatic modes. To perform numerical simulations of PBH formation, one must establish the initial conditions for the spacetime geometry and the fluid in the system. This involves setting the initial profile of curvature fluctuations at the super-horizon scale before the horizon entry to be consistent with the assumption of the growing adiabatic modes. These analytical constructions are typically achieved using the gradient expansion techniques [11, 16, 21].
@@
In a spherically symmetric system, we introduce a type II primordial fluctuation characterized by a non-monotonic behavior of the areal radius $R=\sqrt{A/4\pi}$, where $A$ is the area of the 2-sphere as a function of the radial coordinate $r$ with constant $t$. These fluctuations exhibit distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$. In cosmological perturbation theory, the areal radius $R$ is often expressed via the curvature fluctuation $\zeta$, $R\propto re^{\zeta}$. The type II fluctuation is only realized by a sufficiently large amplitude of $\zeta$ with the scale fixed.
@@
Kopp, Hofmann, and Weller (KHW) [1] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$). They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation. They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold. Carr and Harada (2015) [24] analytically discussed type II PBHs with more general equations of state using simplified models. Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [1, 24]. We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
@@
When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of PBHs originating from them is typically suppressed. However, as mentioned in Ref. [25], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian. Moreover, in the context of a single-field inflationary model, it has been recently discovered in Ref. [26] that type II fluctuations surround the bubbles formed when the inflaton overshoots the barrier. As non-Gaussianities increase, the production of vacuum bubbles becomes dominant, leading to a mass function and abundance primarily driven by these fluctuations. In addition, a similar BH formation mechanism due to the birth of a baby universe with a tunneling process during inflation has been reported in Refs. [27, 28]. The baby universe is connected to our universe through a throat inside the BH, like the neck structure in the type II fluctuation.
@@
If one would like to follow the time evolution of the type II fluctuations via a usual areal radial coordinate system such as in Ref. [21], the neck throat structure would be seriously difficult to deal with. In the Misner–Sharp formalism [29] (see also Refs. [1, 30]), the evolution equations suffer from the trouble associated with $/0$ 0 term at the neck radius satisfying $\partial_{r}R=0$ when we consider type II fluctuations. To overcome this problem, it is necessary to conveniently use and modify the equations, as done in Ref. [27] for another scenario of PBH formation from the collapse of vacuum bubbles.
@@
On the other hand, in this paper, we use the simulation code developed from a 3+1 dimensional simulation code [31] with BSSN [32, 33] formalism with CARTOON method [34], so that we can avoid the same problem.
@@
This paper is organized as follows. Section 2 describes the setup and numerical code of spherically symmetric numerical simulation of PBH formation. In Section 3.1, we present the simulation results, highlighting the time evolution of fluctuations. Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resultant trapping horizon configurations in Section 3.3. In Section 4, we examine the evolution of PBH mass. In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as outlined by KHW. Finally, we offer our conclusions in the last section. Appendix A provides the conformal diagrams of the exact three-zone solutions for PBH formation. Appendix B provides the derivation of the curvature perturbation for the three-zone model. Appendix C proves the appearance of two peaks in the compaction function and the averaged density perturbation for the type II fluctuation.
@@
Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, i.e., $c=G=1$.
@@ 2 Initial data setting
We start our simulation from the time when the length scale of the fluctuation is much larger than the Hubble length $H_{\text{b}}^{-1}$, that is, $k^{-1}\gg(aH_{\text{b}})^{-1}$ with $k$, $a=a(t)$ and $H_{\text{b}}:=(da/dt)/a$ being the characteristic comoving wave number, the scale factor and the Hubble parameter of the background flat FLRW universe, respectively. We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced111 One would use the other expression, $\epsilon:=1/(H_{\rm b}R_{\rm m})$ with $R_{\rm m}\propto a/k$ being the areal radius, where $\mathcal{C}_{\rm SS}(r)$ takes a maximum. , and the zero-th order spatial metric is given by [35]
@@
with $f_{ij}$ being the reference 3-metric of the background flat geometry, and $\partial_{i}\zeta/(aH_{\rm b})=\mathcal{O}(\epsilon)$. We note that the function $\zeta$ is a time-independent arbitrary function of the spatial coordinates $x^{k}$. Once we initially set $\zeta$ as a function of the spatial coordinates $x^{i}$ and fix the gauge conditions, we can successively obtain the growing mode solution up to the following leading terms in this expansion [11, 21]. These long-wavelength solutions are used as the initial data for the time evolution. In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $-3H_{\rm b}$ and the shift vector is equal to zero.
In a spherically symmetric system, the line element at the zero-th order in powers of $\epsilon$ is denoted as
@@
with a parameter $\eta=20$ [31]. In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that around the boundary $z=r=L$;
@@
Up through section 4, we use the following Gaussian-shaped profile in the numerical simulation [31],
@@
where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function [31] introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain.222 The explicit functional form [31] is 11 1$\quad$1 $\text{for}\quad 0\leq r\leq r_{\text{w}},$ $1-\frac{\quantity((r_{\text{w}}-L)^{6}-(L-r)^{6})^{6}}{(r_{\text{w}}-L)^{36}}\quad$ $\text{for}\quad r_{\text{w}}\leq r\leq L,$ (2.6) 00 0$\quad$ $\text{for}\quad L\leq r,$ where $r_{\rm w}$ is set to $0.8L$.
Figure 1 shows the functional form of this profile (left) and the areal radius $R$ (right). By taking a sufficiently large amplitude of the fluctuation $\mu$, the areal radius $R(r)$ of the perturbed region becomes non-monotonic along the radial coordinate $r$, i.e., a pair of the local minimum and maximum points appears with $\partial_{r}R=0$ for $\mu\gtrsim 1.4$ as in Fig. 1. We call this structure neck or throat structure, and KHW defined such fluctuations as of type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
Figure 1: Initial profile of the curvature fluctuation $\zeta$ (left) and areal radius $R$ as functions of the scale-up coordinate radius $z/L$. By taking the amplitude parameter $\mu$ larger, a type II fluctuation appears for $\mu=1.4$ and above with the neck $\partial_{r}R=0$.
The feature of type II fluctuations can also be understood as two peaks333 It is shown in Ref. [26] that the appearance of fluctuations of type II naturally appears via a specific single-field inflation model with a bump, but in this case, one of the peaks of $\mathcal{C}_{\rm SS}$ is hidden inside the bubble. of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ defined by
@@
This function has been firstly introduced in Ref. [11] as an estimator for the criterion of PBH formation, and its usefulness has been recognized through subsequent related works (see Ref. [36] and references therein for details including its relations to legitimate compaction functions in different gauges and physical interpretations). The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for the PBH formation. We will give a general proof for the existence of two peaks and the minimum between them for the compaction function in the type II fluctuation in Appendix C. We focus on the two peak configurations associated with type II fluctuations (see also Ref. [37] for two peak configurations with fluctuations of type I ). The form of the compaction function for the specific function of $\zeta$ (2.5) is shown in Fig. 2. By using the position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threshold value $\mu_{\rm c,II}=e/2=1.359\cdots$ for fluctuations of type II.444 Here, we ignore the window function $W(r)$ to estimate the location $r_{\text{m}}=\sqrt{2}/k$ of the peak of compaction function $\mathcal{C}_{\rm SS}(r)$ since this acts on $\zeta(r)$ only near the outer boundary, $r=r_{\rm w}=0.8L$ For $\mu>\mu_{\rm c,II}$ the compaction function takes the maximum value of $0.5$ at least one location with the radial coordinate, and the number of peak positions becomes two, $r_{m1}$ and $r_{m2}$ with $r_{m1}<r_{m2}$. This implies that the type II fluctuations always reach the threshold for collapse and form PBHs irrespectively on the initial profile of $\zeta(r)$. The isotropic coordinate $r$ does not always give a geometrical size of the enclosed region and may often lead to misunderstanding. Therefore, we also explicitly show the compaction function $\mathcal{C}_{\rm SS}$ as a function of the areal radius $R$ in the right panel of Fig. 2, which will be contrasted with the right panel of Fig. 13.
Figure 2: Compaction function $\mathcal{C}_{\text{SS}}$. The horizontal axes are the scale-up coordinate radius $z/L$ (left) and the areal radius $R$ (right). By taking the amplitude $\mu$ larger, the number of peaks becomes two for $\mu\simeq 1.4$ and above, corresponding to type II fluctuations.
@@
To follow the time evolution of gravitational collapse from the initial profile of fluctuation, we use the numerical code COSMOS-S [38, 39, 40] based on a 3+1 dimensional code of fourth order Runge–Kutta method with BSSN formalism [32, 33]. This code is specialized for spherically symmetric systems using the CARTOON method [34]. We consider the time evolution of the initial fluctuation given by the functional form of $\zeta$ (2.5) with the equation of state parameter $w=1/3$, namely, radiation fluid. Here, the asymptotically spatially flat FLRW boundary condition is implemented as in Ref. [11]. However, we could not avoid the constraint violation at the outer boundary, and this propagates into the bulk region. Therefore, in the following analyses, we terminate the time evolution before the violation reaches the central region relevant for the analyses of PBH formation.
@@
We also find the constraint violation near the center for cases where an apparent horizon forms. Nevertheless, since we find both independent future-directed null directions are inward there, the constraint violation does not propagate outward. In practice, to avoid a crash of numerical calculation, we excise several grids inside the apparent horizon if needed.
@@
We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or, equivalently, $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity. We fix the ratio between the background initial Hubble length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ which ensure the long-wavelength approximation. For a longer duration of the time evolution, we need to take a larger region than the initial Hubble length to avoid the propagation of the constraint violation. In the following, we take $H_{\text{b}}^{-1}=0.01L$ or $H_{\text{b}}^{-1}=0.005L$ depending on the duration of the required time evolution. The horizon entry time $t_{H}$ is defined as $k^{-1}=(aH_{\text{b}})^{-1}|_{t=t_{H}}$ throughout this paper.
3 Spacetime structure and a new definition of type I/II PBHs

### v6

@@ Abstract
This study investigates the formation of primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe. We find that, for a sufficiently large initial amplitude, the configuration of trapping horizons shows characteristic structure due to the existence of bifurcating trapping horizons. We call this type of configuration of the trapping horizons type B PBH, while the structure without a bifurcating trapping horizon type A PBH. As shown in Ref. [PhysRevD.83.124025], in the matter-dominated universe, the type B PBH can be realized by the type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case, type A PBH with a monotonic areal radius (type I fluctuation). Our research reveals that a type II fluctuation does not necessarily result in a type B PBH in the radiation-dominated case. We also find that for an initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with increasing the initial amplitude, depending on its specific profile rather than its fluctuation type.
@@ 1 Introduction
Primordial black holes (PBHs) play a crucial role in exploring the early universe, particularly in investigating inhomogeneities whose scales are much smaller than those of the fluctuations probed by the cosmological microwave background. Since PBHs are remnants of nonlinear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity. They are considered as one of the candidates for dark matter and can be detected as black hole (BH) binaries by gravitational wave interferometers. According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [Zeldovich:1967lct, 10.1093/mnras/152.1.75, 10.1093/mnras/168.2.399, 1978SvA....22..129N, 1980AZh....57..250N].
@@
The mechanism of PBH formation can be briefly understood by considering a model of the gravitational collapse in the flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe (see, e.g., Refs. [PhysRevD.88.084051, galaxies10060112]). A criterion for the collapse has been first proposed by Carr [osti_4109902] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively. A more accurate threshold value has been explored by numerically analyzing general relativistic nonlinear dynamics [Niemeyer:1999ak, PhysRevD.60.084002, Musco_2005, PhysRevD.88.084051, Nakama_2014, Nakama_2014double, osti_4109902, Polnarev_2007, Musco:2008hv, Musco:2012au, ESCRIVA2020100466, Escriva:2019phb, escriva_2021analy, Harada:2015yda, Musco:2018rwt, Musco:2020jjb].
@@
In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$, which represents the growing adiabatic modes. To perform numerical simulations of PBH formation, one must establish the initial conditions for the spacetime geometry and the fluid in the system. This involves setting the initial profile of curvature fluctuations at the super-horizon scale before the horizon entry to be consistent with the assumption of the growing adiabatic modes. These analytical constructions are typically achieved using the gradient expansion techniques [PhysRevD.60.084002, Polnarev_2007, Harada:2015yda].
@@
In a spherically symmetric system, we introduce a type II primordial fluctuation characterized by a non-monotonic behavior of the areal radius $R=\sqrt{A/4\pi}$, where $A$ is the area of the 2-sphere as a function of the radial coordinate $r$ with constant $t$. These fluctuations exhibit distinct neck-like structures, with a minimal areal radius at $\partial_{r}R=0$. In cosmological perturbation theory, the areal radius $R$ is often expressed via the curvature fluctuation $\zeta$, $R\propto re^{\zeta}$. The type II fluctuation is only realized by a sufficiently large amplitude of $\zeta$ with the scale fixed.
@@
Kopp, Hofmann, and Weller (KHW) [PhysRevD.83.124025] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$). They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation. They explicitly showed that it describes a spacetime of BH formation, which they call type II PBH. In addition, they revealed distinct properties of the type II PBH from the type I PBH, which arises from typical type I fluctuations with the amplitude above the threshold. Carr and Harada [PhysRevD.91.084048] analytically discussed type II PBHs with more general equations of state using simplified models. Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [PhysRevD.83.124025, PhysRevD.91.084048]. We aim to investigate the dynamics and the spacetime structure of type II PBH formation in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
@@
When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are statistically unlikely to be generated, and the abundance of type II PBHs originating from them is typically suppressed. However, as mentioned in [Gow:2022jfb], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian. Moreover, in the context of a single-field inflationary model, it has been recently discovered in Ref. [escriva2023formation] that type II fluctuations surround the bubbles formed when the inflaton overshoots the barrier. As non-Gaussianities increase, the production of vacuum bubbles becomes dominant, leading to a mass function and abundance primarily driven by these fluctuations. In addition, a similar BH formation mechanism due to the birth of a baby universe with a tunneling process during inflation has been reported in Refs. [Deng_2017, Deng_2017b]. The baby universe is connected to our universe through a throat inside the BH, like the neck structure in the type II fluctuation.
@@
For example, in the Misner–Sharp formalism [PhysRev.136.B571] (see also Refs. [PhysRevD.83.124025, PhysRevD.74.084013]), the time evolution of the type II fluctuations via a usual areal radial coordinate system such as in Ref. [Harada:2015yda], the neck throat structure would be complicated. The evolution equations suffer from the trouble associated with $0/0$ term at the neck radius satisfying $\partial_{r}R=0$ when we consider type II fluctuations. To overcome this problem, it will be necessary to conveniently modify the equations, as done in Ref. [Deng_2017] for another scenario of PBH formation from the collapse of vacuum bubbles.
@@
In this paper, we use the simulation code developed from a 3+1 dimensional simulation code [Yoo_2019] with BSSN [shibata1995evolution, PhysRevD.59.024007] formalism with CARTOON method [doi:10.1142/S0218271801000834] so that we can avoid the same problem.
@@
This paper is organized as follows. Section 2 describes the setup and numerical code of spherically symmetric numerical simulation of PBH formation. In Sec. 3.1, we present the simulation results, highlighting the time evolution of fluctuations. Then, we explore the formation of trapping horizons, including the BH horizon in Sec. 3.2, and give an interpretation of the resulting trapping horizon configurations in Sec. 3.3. In Sec. 4, we examine the evolution of PBH mass. In Sec. 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as described by KHW. Finally, we offer our conclusions in the last section. Appendix A provides the conformal diagrams of the exact three-zone solutions for PBH formation. Appendix B provides the derivation of the curvature perturbation for the three-zone model. Appendix C proves the appearance of two peaks in the compaction function and the averaged density perturbation for the type II fluctuation.
@@
Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, that is, $c=G=1$.
@@ 2 Initial data setting
We start our simulation from the time when the length scale of the fluctuation is much larger than the Hubble length $H_{\text{b}}^{-1}$, that is, $k^{-1}\gg(aH_{\text{b}})^{-1}$ with $k$, $a=a(t)$ and $H_{\text{b}}:=(da/dt)/a$ being the characteristic comoving wave number, the scale factor and the Hubble parameter of the background flat FLRW universe, respectively. We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced [^1] One would use the other expression, $\epsilon:=1/(H_{\rm b}R_{\rm m})$ with $R_{\rm m}\propto a/k$ being the areal radius, where $\mathcal{C}_{\rm SS}(r)$ takes a maximum. , and the zeroth order spatial metric is given by [Lyth_2005]
@@
with $f_{ij}$ being the reference 3-metric of the background flat geometry, and $\partial_{i}\zeta/(aH_{\rm b})=\mathcal{O}(\epsilon)$. We note that the function $\zeta$ is a time-independent arbitrary function of the spatial coordinates $x^{k}$. Once we initially set $\zeta$ as a function of the spatial coordinates $x^{i}$ and fix the gauge conditions, we can successively obtain the growing mode solution up to the following leading terms in this expansion [PhysRevD.60.084002, Harada:2015yda]. These long-wavelength solutions are used as the initial data for the time evolution. In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $3H_{\rm b}$ and the shift vector is equal to zero.
In a spherically symmetric system, the line element at the zeroth order in powers of $\epsilon$ is denoted as
@@
with a parameter $\eta=20$ [Yoo_2019]. In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that near the boundary $z=r=L$;
@@
Up through Sec. 4, we use the following Gaussian-shaped profile in the numerical simulation [Yoo_2019],
@@
where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function [Yoo_2019] introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain. [^2] The explicit functional form [Yoo_2019] is $1\quad$ $\text{for}\quad 0\leq r\leq r_{\text{w}},$ $1-\frac{\quantity((r_{\text{w}}-L)^{6}-(L-r)^{6})^{6}}{(r_{\text{w}}-L)^{36}}\quad$ $\text{for}\quad r_{\text{w}}\leq r\leq L,$ (2.6) $0\quad$ $\text{for}\quad L\leq r,$ where $r_{\rm w}$ is set to $0.8L$.
Figure 1 shows the functional form of this profile (left) and the areal radius $R$ (right). By taking a sufficiently large amplitude of the fluctuation $\mu$, the areal radius $R(r)$ in the overdense region becomes non-monotonic along the radial coordinate $r$, i.e., a pair of the local minimum and maximum points appears with $\partial_{r}R=0$ for $\mu\gtrsim 1.4$ as in Fig. 1. We call this structure the neck or throat structure, and KHW defined such fluctuations as type II and distinguished them from type I fluctuations, which have a monotonic areal radius [PhysRevD.83.124025].
Figure 1: In the left panel, the vertical and horizontal axes are the functional forms for the curvature perturbation $\zeta(r)$ given by Eq. (2.5) and the scale-up coordinate $z/L$ defined in Eq. (2.3), respectively. Note that $z$ is a monotonically increasing function of $r$ and that $L$ is chosen for the computation region’s length. We plot $\zeta(r)$ for the set of the amplitude parameter values $\mu=0.5,1.2,1.4,1.6$ and $1.8$. In the right panel, the vertical and horizontal axes are the areal radius $R(r)$ and the scale-up coordinate $z/L$, respectively, for the initial data sets of the long-wavelength solutions generated by $\zeta(r)$ plotted in the left panel. We can see that $R(r)$ is a monotonically increasing function for $\mu=0.5$ and $1.2$, while it increases, decreases, and increases again as $r$ increases for $\mu=1.4,1.6$, and $1.8$. In the latter cases, $R(r)$ has a maximum and a minimum with $\partial_{r}R=0$, where the minimum corresponds to the neck.
The feature of type II fluctuations can also be understood as two peaks [^3] It is shown in Ref. [escriva2023formation] that the appearance of fluctuations of type II naturally appears via a specific single-field inflation model with a bump, but in this case, one of the peaks of $\mathcal{C}_{\rm SS}$ is hidden inside the bubble. of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ at super horizon scales, which is calculated as [Harada:2015yda]
@@
The Shibata-Sasaki compaction function was first introduced in Ref. [PhysRevD.60.084002] as an estimator for the criterion of PBH formation. See Refs. [Harada:2023ffo, Harada:2024trx] and references therein for the initial confusion in its introduction, its usefulness, its relation to legitimate compaction functions in different gauges, and its geometrical origin as a conformal compactness function. The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for PBH formation. We will give a general proof for two peaks and the minimum between them for the compaction function in the type II fluctuation in the Appendix C. We focus on the two peak configurations associated with type II fluctuations (see also Ref. [Escriva:2023qnq] for two peak configurations with fluctuations of type I ). The form of the compaction function for the specific function of $\zeta$ (2.5) is shown in Fig. 2. By using the position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threshold value $\mu_{\rm c,II}=e/2=1.359\cdots$ for fluctuations of type II. [^4] Here, we ignore the window function $W(r)$ to estimate the location $r_{\text{m}}=\sqrt{2}/k$ of the peak of compaction function $\mathcal{C}_{\rm SS}(r)$ since this acts on $\zeta(r)$ only near the outer boundary, $r=r_{\rm w}=0.8L$ . For $\mu>\mu_{\rm c,II}$ the compaction function takes the maximum value of $0.5$ at least one location with the radial coordinate, and the number of peak positions becomes two, $r_{m1}$ and $r_{m2}$ with $r_{m1}<r_{m2}$. This implies that type II fluctuations always reach the collapse threshold and form PBHs irrespectively of the initial $\zeta(r)$ profile. The isotropic coordinate $r$ does not always give the geometrical size of the enclosed region, which can often lead to misunderstanding. Therefore, we also explicitly show the compaction function $\mathcal{C}_{\rm SS}$ as a function of the areal radius $R$ in the right panel of Fig. 2.
Figure 2: In the left panel, the vertical and horizontal axes are the compaction function $\mathcal{C}_{\text{SS}}(r)$ and the scale-up coordinate $z/L$, respectively, for the initial data sets of the long-wavelength solutions generated by Eq. (2.5) for the same set of the amplitude parameter $\mu$. We can see that $\mathcal{C}_{\text{SS}}(r)$ takes only one maximum, which is smaller than $1/2$, for $\mu=0.5$ and $1.2$, while it takes two distinct maxima, which are equal to $1/2$, and a single minimum between the two maxima for $\tilde{\mu}=1.4,1.6$ and $1.8$. The latter cases correspond to type II fluctuations as shown in Appendix C. In the right panel, the vertical and horizontal axes are the compaction function $\mathcal{C}_{\text{SS}}(r)$ and the areal radius $R(r)$, respectively, for the same initial data sets. We can see that $\mathcal{C}_{\text{SS}}(r)$ takes only the maximum or the two maxima at the larger areal radius $R$ for the larger values of $\mu$. This implies that the mass of the formed black hole increases as $\mu$ increases, as seen in Fig. 12.
@@
To follow the time evolution of gravitational collapse from the initial profile of fluctuation, we use the numerical code COSMOS-S [PhysRevLett.111.161102, PhysRevD.89.104032, PhysRevD.105.103538] based on a 3+1 dimensional code of fourth order Runge–Kutta method with BSSN formalism [shibata1995evolution, PhysRevD.59.024007]. This code is specialized for spherically symmetric systems using the CARTOON method [doi:10.1142/S0218271801000834]. We consider the time evolution of the initial fluctuation given by the functional form of $\zeta$ (2.5) with the equation of state parameter $w=1/3$, namely, radiation fluid. Here, the asymptotically spatially flat FLRW boundary condition is implemented as in Ref. [PhysRevD.60.084002]. However, we could not avoid violating the constraint at the outer boundary, which propagates into the bulk region. Therefore, we terminate the time evolution before the violation reaches the central region relevant to the PBH formation in the simulation.
@@
We also find the constraint violation near the center for cases where an apparent horizon forms. Nevertheless, since we find both independent future-directed null directions inward, the constraint violation does not propagate outward. In practice, we excise several grids inside the apparent horizon to avoid the numerical calculation from crashing if needed.
@@
We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or equivalently $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity. We fix the ratio between the initial Hubble background length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ that ensures the long-wavelength approximation. To avoid propagating the constraint violation, we need to take a larger region than the initial Hubble length for a longer duration of the time evolution. In the following, we take $H_{\text{b}}^{-1}=0.01L$ or $H_{\text{b}}^{-1}=0.005L$ depending on the duration of the required time evolution. The horizon entry time $t_{H}$ is defined as $k^{-1}=(aH_{\text{b}})^{-1}|_{t=t_{H}}$ throughout this paper.
3 Spacetime structure and definition of type A/B PBHs